If the internal policy of a department store is to tolerate shortages up to 4% (meaning the service level is 96%) of any given products on sale. How many items of a given product should be stocked, knowing that the sale of this product is normally distributed with a mean of 85 units and a standard deviation of 5 units?
Answer #1

Given that,
Margin of error= E= 4%=0.04
σ= 5
(1–α)%=96%
(1–α)=0.96
α=0.04
α/2=0.02
Zα/2=2.05 ....... from standard normal table.
Sample size = n={(Zα/2 ×σ)/E}^2
={(2.05×5)//0.04}^2   
=65664.0625
= 65664... nearest whole number.